The rms velocity of hydrogen is √7 times t5he rms velocity of nitrogen. If T is the temperature of the gas
a. T(H2) = T(N2)
b. T(H2) > T(N2)
c. T(H2) < T(N2)
d. T(H2) = √7 T(N2)

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. Root mean square speed Urms = √(3RT/M) ⇒ (u_rms 〖(H〗_2) )/(u_rms (N_2 ) ) =√7 = √((〖T (H〗_1) )/2 × 28/(T (N_2))) ⇒ 7 = (14T 〖(H〗_2) )/(T (N_2 ) ) ⇒ T (N2) = 2T (H2), ie, T (H2) < T (N2)